OpenStudy (anonymous):

2^(1-x)=3^x Can someone please help me solve this?

OpenStudy (jack1):

use log's

OpenStudy (anonymous):

how do it do it? i know i have to use logs but i am completely lost as of how to start

OpenStudy (jack1):

2^(1-x) = 3^x Multiply both sides by 2^(x-1): 1 = 2^(x-1) 3^x 1 = 2^(x-1) 3^x is equivalent to 2^(x-1) 3^x = 1: 2^(x-1) 3^x = 1 2^(x-1) 3^x = e^(log(2^(x-1))) e^(log(3^x)) = e^((x-1) log(2)) e^(x log(3)) = e^((x-1) log(2)+x log(3)): e^(log(2) (x-1)+log(3) x) = 1 Take the natural logarithm of both sides: ln(2) (x-1)+ln(3) x = 0 (log(2)+log(3)) x-log(2) = 0 =(log(2)+log(3)) x = log(2)

OpenStudy (jack1):

so x = ln2 / (ln2 + ln3)

OpenStudy (anonymous):

Thank you so much!!! you explained it perfectly :)